This is factoring and solving polynomial equations
In this case it would be a trinomial
Simplify all square roots and answers should not be decimals
Factor or solve the following:
X⁴-16x²+48

This question might require using the slide rule which I'm not 100% sure if it will but consider using it. Thank you in advance to whoever attempts this prob!

Found 2 solutions by ewatrrr, MathTherapy:
Answer by ewatrrr(24785)


Hi
x^4-16x^2+48
X^2 - 16X + 48 = 0          | Let  X = x^2 
(X - 12)(X - 4) = 0
 X = 12      0r   X = 4
 x^2 = 12         x^2 = 4                
  x = ± 2√3         x = ± 2
Wish You the Best in your Studies.

Answer by MathTherapy(10552)


This is factoring and solving polynomial equations
In this case it would be a trinomial
Simplify all square roots and answers should not be decimals
Factor or solve the following:
X⁴-16x²+48

This question might require using the slide rule which I'm not 100% sure if it will but consider using it. Thank you in advance to whoever attempts this prob!
This needs to be FACTORED since there's no EQUALS sign, as: 
I have absolutely no IDEA why the other person solved this when it's extremely clear that it's NOT an equation!